# PACE: Procedural Abstractions for Communicating Efficiently

### v3

@@ Introduction
Procedural tasks such as cooking and programming require executing a sequence of actions to achieve a desired goal. A natural approach to reduce their complexity and improve generalisation to new tasks is to introduce abstractions for common sequences of actions [Solway et al., 2014]. For example, in cooking, techniques such as sautéing or kneading serve as foundational building blocks that simplify complex recipes. Similarly, abstractions emerge in repeated communication between human dyads when collaborating on shared tasks [Krauss and Weinheimer, 1964, Hawkins et al., 2020, McCarthy et al., 2021]. Over time, as new abstractions are introduced into the shared language, communication becomes more concise. This can improve cooperation and allow to reach the goal more easily.
@@
How communication shapes abstractions is of much interest within the AI and Cognitive Science communities [Lee, 1996, Ho et al., 2019]. A compelling principle is Efficient Communication, which argues that languages are under pressure to be informative whilst minimising cognitive load [Kemp and Regier, 2012, Gibson et al., 2017, Zaslavsky et al., 2019, Gibson et al., 2019]. This provides insights into the abstractions humans converge to in semantic domains such as colour, kinships and others [Xu et al., 2020, Regier et al., 2015, Kemp and Regier, 2012, Yin et al., 2024]. Here we want to explore the principle of Efficient Communication with artificial agents in collaborative procedural tasks: how do the pressures for efficient communication manifest and impact abstraction introduction and use?
@@
The architect-builder game introduced in McCarthy et al. [2021] provides a simple framework to study the use of abstractions in collaborative tasks. It is a repeated game with two participants, the architect and the builder, who must complete a collaborative building task. In each round, the architect observes the goal-scene on their screen and conveys some instructions in natural language to the builder. The builder, who cannot see the original scene, interprets these instructions and attempts to reconstruct the shapes by placing blocks within the grid. The architect may send further instructions to help the builder. McCarthy et al. [2021] showed that in early rounds human participants use lengthy utterances, while later realizing it is beneficial to introduce abstractions for commonly occurring shapes (e.g. an L-shape). By doing this they move towards a more compact language, thereby leading to more efficient communication. Creating such abstractions is a key feature of human collaboration. In this work, we develop a computational framework of artificial agents to solve this kind of collaborative task (Figure 1). Previous work provides computational models of this process [McCarthy et al., 2021, Jergéus et al., 2022], but rely on simplifying assumptions, limiting the ability to capture the collaborative dynamics of language learning, which we address in this work.
@@
We propose a novel multi-agent neuro-symbolic method called Procedural Abstractions for Communicating Efficiently (PACE). We integrate both neural and symbolic learning methods from the computer science literature. On the symbolic side, library learning [Ellis et al., 2021, Bowers et al., 2023], a method from program synthesis that aims to abstract common subprograms into new, more easily reusable terms. On the neural side, emergent communication (EC) [Foerster et al., 2016, Lazaridou and Baroni, 2020], and reinforcement learning (RL) [Sutton and Barto, 1998], enable the development of a flexible, learnable communication language. PACE offers a unified framework for studying the formation and evolution of abstractions across multi-round interactions.
@@
In previous work, EC has been used to explain how communicative pressures for efficient communication shape the language structure in other settings, such as Carlsson et al. [2021, 2024]. Here, we apply EC techniques to study how pressures manifest within abstraction learning.
@@
In the Architect-Builder Game, the architect is provided with a set of goal-scenes depicting two adjacent shapes on a (9x9) grid. Each shape is a variable-size combination of 2x1 horizontal and 1x2 vertical blocks. The architect needs to communicate instructions to the builder (who does not know the goal state) that allow it to construct the goal-scene starting from a blank grid. This is inspired by the human experiments in McCarthy et al. [2021].
@@
We extend the dataset from McCarthy et al. [2021] for our experiments. We increase the size going from $3$ to $31$ unique shapes, with multiple sub-shapes reoccurring in different shapes. Our shapes are of different sizes and resemble either uppercase or lowercase letters from the English alphabet. As before, the dataset consists of scenes composed of two shapes placed side by side. Our dataset contains $961$ goal-scenes (compared to the original $9$), which enables us to use it for training neural agents. We split the dataset into 930 training scenes and 31 test scenes. These splits are constructed to ensure the distribution over shapes is the same. However, in the test set, the ordered pair of shapes constituting each goal-scene, does not appear in the training set.
@@
The architect initially has one program to construct each scene from an empty grid. As abstractions are introduced, there will be multiple alternative programs of different lengths available to the architect. They may differ in reconstruction accuracy, in particular, programs containing a new abstraction will initially have lower accuracy, as the builder has not been exposed to them. Hence, this is a classic exploration vs. exploitation problem for which reinforcement learning with bandit techniques222an analogy coming from a gambler repeatedly having to select which arm of a slot machine to play. are well-suited [Sutton and Barto, 1998, Lattimore and Szepesvári, 2020] More specifically, we view the selection between programs as a contextual multi-armed bandit with combinatorial actions: in general each goal-scene has multiple programs to choose from, corresponding to the arms of the bandit. To estimate the quality of each program, we maintain a table of Q-values, one for each action in $\mathcal{A}$, with the quality of the program being the product of the Q-values of its actions: $Q(p)=\prod_{i=1}^{|p|}\gamma Q(a_{i})$. [^3] We empirically found that initialising the Q-values to 00 works best in practice. This captures the trade-off between program length and the communicative accuracy of instructions. We empirically determine a value of $\gamma=0.99$, which results in a small bias for shorter programs. To ensure that the architect also explores new programs, we adopt an $\epsilon$-greedy strategy, where with probability $1-\epsilon$ a random program (arm) is selected. We fix $\epsilon$ at a constant value of $0.1$, ensuring a constant level of exploration. After the communication of the program (see below) we update the estimated value of its component actions as $Q(a)\leftarrow Q(a)+\alpha(r-Q(a))$, where $r$ is the reconstruction accuracy which is 1 if this instruction was successfully interpreted by the builder and 0 otherwise. This approach ensures that the new programs introduced after each Abstraction phase will be explored by the architect.
@@
Having selected a program, $p$, the architect and builder play a one-step signalling game for each of its instructions $a_{i}$. This takes the form of $(x_{i},a_{i},x_{i+1})$: the grid state $x_{i}$ is transformed into $x_{i+1}$ by action $a_{i}$. The architect learns a neural communication policy $\pi_{comm}$, which produces a message $m_{i}=\pi_{comm}(a_{i})$. Similarly, the builder learns a policy $\pi_{bldr}$ which estimates the next grid-state, and is defined as $\widehat{x}_{i+1}=\pi_{bldr}(x_{i},m_{i})$. These are both implemented as fully-connected neural networks. Since our messages are discrete, we use the gumbel-softmax relaxation to sample from discrete messages which makes our model end-to-end differentiable [Jang et al., 2017]. The policies are jointly trained to minimise the binary cross-entropy loss between the target next state $x_{i+1}$ and the builder’s output $\widehat{x}_{i+1}$. We also introduce a bias for positive signalling Eccles et al. [2019] in the architect’s loss to encourage the architect’s messages to carry meaningful information about the instruction they represent.
@@
After some rounds of communication, the architect enters the abstraction phase, where it searches for novel abstractions allowing for shorter programs for describing goal scenes. Abstractions are constructed by an improved version of the procedure used in McCarthy et al. [2021], where we are able to remove the explicit upper-size limit on the library size.
@@
[…]
@@
As in other settings like recursive numeral systems [Denić and Szymanik, 2024], programs in the Architect-Builder game can be perfectly informative even with a minimal vocabulary – precisely of size two, with one term referring to the horizontal primitive block and the other referring to the vertical one. With just the two initial action-words the architect can describe infinitely complex goal-scenes, given enough time! This is because the semantics of the environment are compositional. Thus, as Denić and Szymanik [2024] show for recursive numeral systems, we expect that the pressures in competition to shape the agents’ languages in the Architect-Builder game are the average morphosyntactic complexity of the programs describing the goal-scenes and the language size. In our case this corresponds to a trade-off between the average length of the programs communicated by the architect, and the number of lexicalised terms, i.e. the number of primitives plus abstractions in $\mathcal{A}$, the architect’s lexicon.
@@
The neural agents are deep neural networks which have $1$ and $2$ hidden layers of size $200$, respectively. The architect’s output layer is of size $30$, its maximum vocabulary size. While the output layer of the builder is of size $81$, the size of the grid. We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. We use a learning rate of $0.0009$ in conjunction with the ADAM optimiser [Kingma and Ba, 2014]. The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in the architects symbolic table to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.
@@
PACE does indeed reduce morphosyntactic complexity as seen in Figure 3 (left). Starting from an average morphosyntactic complexity of $9.95$, over successive interaction steps, PACE is able to reduce average morphosyntactic complexity consistently and finally converges around a value of $4.92\pm 0.20$ When we compare PACE to the No abstraction variant, the average morphosyntactic complexity is halved, going from an average of $10$, indicating that a more efficient language has been derived. The Greedy variant reduces program lengths even more, however, this comes at a price. In Figure 3 (right) we see drastic drops in accuracy for Greedy every time an abstraction is introduced, whereas PACE’s slower introduction leads to smaller drops in accuracy. For a while, both strategies manage to recover near full accuracy, but after some time, the Greedy communication fails to recover and becomes progressively worse.
@@
As more terms are introduced and the average morphosyntactic complexity reduces, we find that it becomes harder to introduce new abstract terms. This explains why the Greedy approach diverges as it is forced to use new abstractions in its programs. This raises a profound question which is of interest within efficient communication: why does PACE introduce some abstractions whereas others are omitted? To understand this we conduct an experiment where at each step, before the communicative phase, we gather all candidate abstractions, group them for their size (as in how big the shape the abstraction refers to is in terms of primitive blocks) and frequency in our dataset, and then introduce them one by one into $\mathcal{A}$. For each candidate abstraction, we let the architect and builder communicate until convergence and then probe the final language to check if the tested abstraction has been adopted or not.  [^4] Note that to make this experiment computationally tractable we limit to sampling $3$ abstractions for each frequency and size, wherever we have more than $3$. This provides an exhaustive search of differing abstractions and allows us to analyse their ease of adoption or learnability into the current language [Steinert-Threlkeld and Szymanik, 2020], based on how much reduction in average morphosyntactic complexity they can potentially lead to. At the next step, the experiment is repeated with the new symbolic language $\mathcal{A}$ augmented with the new abstraction introduced at the previous Abstraction phase, making each step of this experiment conditioned on the current language of PACE.
We plot the results in Figure 4. We find that when $\mathcal{A}$ is only composed of primitive terms (the language is small), abstractions are generally always adopted but as $\mathcal{A}$ grows abstractions which would result in languages that are further away from the Pareto Frontier become harder to learn. Eventually, introducing new abstractions becomes infeasible, resulting in PACE’s convergence. This demonstrates how communicative efficiency impact language formation within PACE, and connects with Steinert-Threlkeld and Szymanik [2020] which show that simplicity and ease of learning are intertwined.
@@
McCarthy et al. [2021] propose a Bayesian model of procedural abstractions that combines library learning [Ellis et al., 2021] with social reasoning-based communication [Goodman and Frank, 2016]. While effective in capturing human-like trends, this model assumes a predefined mapping from instructions to builder’s actions, essentially defining a priori the meaning of messages that might not have been introduced yet. They also explicitly constrain language size to prevent adding further abstractions after a certain number have been introduced. Jergéus et al. [2022] take a Deep Reinforcement Learning approach but oversimplify interactions by assuming the builder immediately understands abstractions, collapsing the setup into a single-agent framework and removing pressures for efficient communication. By using EC within PACE, our approach enables both agents to learn interactively, introducing abstractions organically and letting communicative pressures naturally limit language size.
Emergent Communication has been widely used to study various human language phenomena, such as compositionality [Mordatch and Abbeel, 2018, Chaabouni et al., 2020, Ren et al., 2020], the impact of populations [Kim and Oh, 2021, Chaabouni et al., 2022, Rita et al., 2022], and naming conventions in semantic categories [Carlsson et al., 2024]. The architect-builder game presents a very challenging setting, being a repeated game with multiple interactions per round and sparse rewards. By incorporating symbolic methods with emergent communication, we provide a more interpretable, easy-to-use approach for modelling the introduction and use of procedural abstractions into conversational AI dyads.

### v4

@@ Introduction
Procedural tasks such as cooking and programming require executing a sequence of actions to achieve a desired goal. A natural approach to reduce their complexity and improve generalisation to new tasks is to introduce abstractions for common sequences of actions \@BBOPcitep\@BAP\@BBN(Solway et al., 2014)\@BBCP. For example, in cooking, techniques such as sautéing or kneading serve as foundational building blocks that simplify complex recipes. Similarly, abstractions emerge in repeated communication between human dyads when collaborating on shared tasks \@BBOPcitep\@BAP\@BBN(Krauss and Weinheimer, 1964; Hawkins et al., 2020; McCarthy et al., 2021)\@BBCP. Over time, as new abstractions are introduced into the shared language, communication becomes more concise. This can improve cooperation and allow to reach the goal more easily.
@@
How communication shapes abstractions is of much interest within the AI and Cognitive Science communities \@BBOPcitep\@BAP\@BBN(Lee, 1996; Ho et al., 2019)\@BBCP. A compelling principle is Efficient Communication, which argues that languages are under pressure to be informative whilst minimising cognitive load \@BBOPcitep\@BAP\@BBN(Kemp and Regier, 2012; Gibson et al., 2017; Zaslavsky et al., 2019; Gibson et al., 2019)\@BBCP. This provides insights into the abstractions humans converge to in semantic domains such as colour, kinships and others \@BBOPcitep\@BAP\@BBN(Xu et al., 2020; Regier et al., 2015; Kemp and Regier, 2012; Yin et al., 2024)\@BBCP. Here we want to explore the principle of Efficient Communication with artificial agents in collaborative procedural tasks: how do the pressures for efficient communication manifest and impact abstraction introduction and use?
@@
The architect-builder game introduced in \@BBOPcite\@BAP\@BBNMcCarthy et al. (2021)\@BBCP provides a simple framework to study the use of abstractions in collaborative tasks. It is a repeated game with two participants, the architect and the builder, who must complete a collaborative building task. In each round, the architect observes the goal-scene on their screen and conveys some instructions in natural language to the builder. The builder, who cannot see the original scene, interprets these instructions and attempts to reconstruct the shapes by placing blocks within the grid. The architect may send further instructions to help the builder. \@BBOPcitet\@BAP\@BBNMcCarthy et al. (2021)\@BBCP showed that in early rounds human participants use lengthy utterances, while later realizing it is beneficial to introduce abstractions for commonly occurring shapes (e.g. an L-shape). By doing this they move towards a more compact language, thereby leading to more efficient communication. Creating such abstractions is a key feature of human collaboration. In this work, we develop a computational framework of artificial agents to solve this kind of collaborative task (Figure 1). Previous work provides computational models of this process \@BBOPcitep\@BAP\@BBN(McCarthy et al., 2021; Jergéus et al., 2022)\@BBCP, but rely on simplifying assumptions, limiting the ability to capture the collaborative dynamics of language learning, which we address in this work.
@@
We propose a novel multi-agent neuro-symbolic method called Procedural Abstractions for Communicating Efficiently (PACE). We integrate both neural and symbolic learning methods from the computer science literature. On the symbolic side, library learning \@BBOPcitep\@BAP\@BBN(Ellis et al., 2021; Bowers et al., 2023)\@BBCP, a method from program synthesis that aims to abstract common subprograms into new, more easily reusable terms. On the neural side, emergent communication (EC) \@BBOPcitep\@BAP\@BBN(Foerster et al., 2016; Lazaridou and Baroni, 2020)\@BBCP, and reinforcement learning (RL) \@BBOPcitep\@BAP\@BBN(Sutton and Barto, 1998)\@BBCP, enable the development of a flexible, learnable communication language. PACE offers a unified framework for studying the formation and evolution of abstractions across multi-round interactions.
@@
In previous work, EC has been used to explain how communicative pressures for efficient communication shape the language structure in other settings, such as \@BBOPcitet\@BAP\@BBNCarlsson et al. (2021, 2024)\@BBCP. Here, we apply EC techniques to study how pressures manifest within abstraction learning.
@@
In the Architect-Builder Game, the architect is provided with a set of goal-scenes depicting two adjacent shapes on a (9x9) grid. Each shape is a variable-size combination of 2x1 horizontal and 1x2 vertical blocks. The architect needs to communicate instructions to the builder (who does not know the goal state) that allow it to construct the goal-scene starting from a blank grid. This is inspired by the human experiments in \@BBOPcite\@BAP\@BBNMcCarthy et al. (2021)\@BBCP.
@@
We extend the dataset from \@BBOPcite\@BAP\@BBNMcCarthy et al. (2021)\@BBCP for our experiments. We increase the size going from $3$ to $31$ unique shapes, with multiple sub-shapes reoccurring in different shapes. Our shapes are of different sizes and resemble either uppercase or lowercase letters from the English alphabet. As before, the dataset consists of scenes composed of two shapes placed side by side. Our dataset contains $961$ goal-scenes (compared to the original $9$), which enables us to use it for training neural agents. We split the dataset into 930 training scenes and 31 test scenes. These splits are constructed to ensure the distribution over shapes is the same. However, in the test set, the ordered pair of shapes constituting each goal-scene, does not appear in the training set.
@@
The architect initially has one program to construct each scene from an empty grid. As abstractions are introduced, there will be multiple alternative programs of different lengths available to the architect. They may differ in reconstruction accuracy, in particular, programs containing a new abstraction will initially have lower accuracy, as the builder has not been exposed to them. Hence, this is a classic exploration vs. exploitation problem for which reinforcement learning with bandit techniques222an analogy coming from a gambler repeatedly having to select which arm of a slot machine to play. are well-suited \@BBOPcitep\@BAP\@BBN(Sutton and Barto, 1998; Lattimore and Szepesvári, 2020)\@BBCP More specifically, we view the selection between programs as a contextual multi-armed bandit with combinatorial actions: in general each goal-scene has multiple programs to choose from, corresponding to the arms of the bandit. To estimate the quality of each program, we maintain a table of Q-values, one for each action in $\mathcal{A}$, with the quality of the program being the product of the Q-values of its actions: $Q(p)=\prod_{i=1}^{|p|}\gamma Q(a_{i})$. [^3] We empirically found that initialising the Q-values to 0 works best in practice. This captures the trade-off between program length and the communicative accuracy of instructions. We empirically determine a value of $\gamma=0.99$, which results in a small bias for shorter programs. To ensure that the architect also explores new programs, we adopt an $\epsilon$-greedy strategy, where with probability $1-\epsilon$ a random program (arm) is selected. We fix $\epsilon$ at a constant value of $0.1$, ensuring a constant level of exploration. After the communication of the program (see below) we update the estimated value of its component actions as $Q(a)\leftarrow Q(a)+\alpha(r-Q(a))$, where $r$ is the reconstruction accuracy which is 1 if this instruction was successfully interpreted by the builder and 0 otherwise. This approach ensures that the new programs introduced after each Abstraction phase will be explored by the architect.
@@
Having selected a program, $p$, the architect and builder play a one-step signalling game for each of its instructions $a_{i}$. This takes the form of $(x_{i},a_{i},x_{i+1})$: the grid state $x_{i}$ is transformed into $x_{i+1}$ by action $a_{i}$. The architect learns a neural communication policy $\pi_{comm}$, which produces a message $m_{i}=\pi_{comm}(a_{i})$. Similarly, the builder learns a policy $\pi_{bldr}$ which estimates the next grid-state, and is defined as $\widehat{x}_{i+1}=\pi_{bldr}(x_{i},m_{i})$. These are both implemented as fully-connected neural networks. Since our messages are discrete, we use the gumbel-softmax relaxation to sample from discrete messages which makes our model end-to-end differentiable \@BBOPcitep\@BAP\@BBN(Jang et al., 2017)\@BBCP. The policies are jointly trained to minimise the binary cross-entropy loss between the target next state $x_{i+1}$ and the builder’s output $\widehat{x}_{i+1}$. We also introduce a bias for positive signalling \@BBOPcite\@BAP\@BBNEccles et al. (2019)\@BBCP in the architect’s loss to encourage the architect’s messages to carry meaningful information about the instruction they represent.
@@
After some rounds of communication, the architect enters the abstraction phase, where it searches for novel abstractions allowing for shorter programs for describing goal scenes. Abstractions are constructed by an improved version of the procedure used in \@BBOPcitet\@BAP\@BBNMcCarthy et al. (2021)\@BBCP, where we are able to remove the explicit upper-size limit on the library size.
@@
where $a_{cand}$ is the new candidate abstraction, $\{p_{i}\}_{i=1}^{N}$ are the known programs so far. The right side of the equation is further defined in 2. It provides a measure of the expected reduction in program length, which is defined in terms of the minimum description length ($\mathit{MDL}$) –the shortest program achievable using also the new candidate abstraction.
$P(p_{i}|\mathcal{A}\cup\{a_{cand}\})=exp(-\mathit{MDL}(p_{i}|\mathcal{A}\cup\{a_{cand}\}))$ (2)
@@
As in other settings like recursive numeral systems \@BBOPcitep\@BAP\@BBN(Denić and Szymanik, 2024)\@BBCP, programs in the Architect-Builder game can be perfectly informative even with a minimal vocabulary – precisely of size two, with one term referring to the horizontal primitive block and the other referring to the vertical one. With just the two initial action-words the architect can describe infinitely complex goal-scenes, given enough time! This is because the semantics of the environment are compositional. Thus, as \@BBOPcite\@BAP\@BBNDenić and Szymanik (2024)\@BBCP show for recursive numeral systems, we expect that the pressures in competition to shape the agents’ languages in the Architect-Builder game are the average morphosyntactic complexity of the programs describing the goal-scenes and the language size. In our case this corresponds to a trade-off between the average length of the programs communicated by the architect, and the number of lexicalised terms, i.e. the number of primitives plus abstractions in $\mathcal{A}$, the architect’s lexicon.
@@
The neural agents are deep neural networks which have $1$ and $2$ hidden layers of size $200$, respectively. The architect’s output layer is of size $30$, its maximum vocabulary size. The output layer of the builder is of size $81$, corresponding to the size of the grid (9x9). We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. We use a learning rate of $0.0009$ in conjunction with the ADAM optimiser \@BBOPcitep\@BAP\@BBN(Kingma and Ba, 2014)\@BBCP. The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in the architects symbolic table to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.
@@
PACE does indeed reduce morphosyntactic complexity as seen in Figure 3 (left). Starting from an average morphosyntactic complexity of $9.95$, over successive interaction steps, PACE is able to reduce average morphosyntactic complexity consistently and finally converges around a value of $4.92\pm 0.20$ . When we compare PACE to the No abstraction variant, the average morphosyntactic complexity is halved, going from an average of $10$, indicating that a more efficient language has been derived. The Greedy variant reduces program lengths even more, however, this comes at a price. In Figure 3 (right) we see drastic drops in accuracy for Greedy every time an abstraction is introduced, whereas PACE’s slower introduction leads to smaller drops in accuracy. For a while, both strategies manage to recover near full accuracy, but after some time, the Greedy communication fails to recover and becomes progressively worse.
@@
As more terms are introduced and the average morphosyntactic complexity reduces, we find that it becomes harder to introduce new abstract terms. This explains why the Greedy approach diverges as it is forced to use new abstractions in its programs. This raises a profound question which is of interest within efficient communication: why does PACE introduce some abstractions whereas others are omitted? To understand this we conduct an experiment where at each step, before the communicative phase, we gather all candidate abstractions, group them for their size (as in how big the shape the abstraction refers to is in terms of primitive blocks) and frequency in our dataset, and then introduce them one by one into $\mathcal{A}$. For each candidate abstraction, we let the architect and builder communicate until convergence and then probe the final language to check if the tested abstraction has been adopted or not.  [^4] Note that to make this experiment computationally tractable we limit to sampling $3$ abstractions for each frequency and size, wherever we have more than $3$. This provides an exhaustive search of differing abstractions and allows us to analyse their ease of adoption or learnability into the current language \@BBOPcitep\@BAP\@BBN(Steinert-Threlkeld and Szymanik, 2020)\@BBCP, based on how much reduction in average morphosyntactic complexity they can potentially lead to. At the next step, the experiment is repeated with the new symbolic language $\mathcal{A}$ augmented with the new abstraction introduced at the previous Abstraction phase, making each step of this experiment conditioned on the current language of PACE.
We plot the results in Figure 4. We find that when $\mathcal{A}$ is only composed of primitive terms (the language is small), abstractions are generally always adopted but as $\mathcal{A}$ grows abstractions which would result in languages that are further away from the Pareto Frontier become harder to learn. Eventually, introducing new abstractions becomes infeasible, resulting in PACE’s convergence. This demonstrates how communicative efficiency impact language formation within PACE, and connects with \@BBOPcitet\@BAP\@BBNSteinert-Threlkeld and Szymanik (2020)\@BBCP which show that simplicity and ease of learning are intertwined.
@@
\@BBOPcitet\@BAP\@BBNMcCarthy et al. (2021)\@BBCP propose a Bayesian model of procedural abstractions that combines library learning \@BBOPcitep\@BAP\@BBN(Ellis et al., 2021)\@BBCP with social reasoning-based communication \@BBOPcitep\@BAP\@BBN(Goodman and Frank, 2016)\@BBCP. While effective in capturing human-like trends, this model assumes a predefined mapping from instructions to builder’s actions, essentially defining a priori the meaning of messages that might not have been introduced yet. They also explicitly constrain language size to prevent adding further abstractions after a certain number have been introduced. \@BBOPcitet\@BAP\@BBNJergéus et al. (2022)\@BBCP take a Deep Reinforcement Learning approach but oversimplify interactions by assuming the builder immediately understands abstractions, collapsing the setup into a single-agent framework and removing pressures for efficient communication. By using EC within PACE, our approach enables both agents to learn interactively, introducing abstractions organically and letting communicative pressures naturally limit language size.
Emergent Communication has been widely used to study various human language phenomena, such as compositionality \@BBOPcitep\@BAP\@BBN(Mordatch and Abbeel, 2018; Chaabouni et al., 2020; Ren et al., 2020)\@BBCP, the impact of populations \@BBOPcitep\@BAP\@BBN(Kim and Oh, 2021; Chaabouni et al., 2022; Rita et al., 2022)\@BBCP, and naming conventions in semantic categories \@BBOPcitep\@BAP\@BBN(Carlsson et al., 2024)\@BBCP. The architect-builder game presents a very challenging setting, being a repeated game with multiple interactions per round and sparse rewards. By incorporating symbolic methods with emergent communication, we provide a more interpretable, easy-to-use approach for modelling the introduction and use of procedural abstractions into conversational AI dyads.